# SmartIntentNN: Towards Smart Contract Intent Detection

### v3

@@ Abstract
Smart contracts on the blockchain offer decentralized financial services but often lack robust security measures, resulting in significant economic losses. Although substantial research has focused on identifying vulnerabilities, a notable gap remains in evaluating the malicious intent behind their development. To address this, we introduce SmartIntentNN (Smart Contract Intent Neural Network), a deep learning-based tool designed to automate the detection of developers’ intent in smart contracts. Our approach integrates a Universal Sentence Encoder for contextual representation of smart contract code, employs a K-means clustering algorithm to highlight intent-related code features, and utilizes a bidirectional LSTM-based multi-label classification network to predict ten distinct types of high-risk intent. Evaluations on a dataset of 10,000 smart contracts demonstrate that SmartIntentNN surpasses all baselines, achieving an F1-score of up to 0.8633.
@@ I Introduction
A smart contract is a type of computer program and transaction protocol, engineered to execute, control, or document legally binding events and actions automatically according to the stipulations of a contract or agreement [1]. Users generally interact with smart contracts by initiating transactions to invoke various functions. From a programming standpoint, current research on smart contract security predominantly focuses on identifying vulnerabilities and defects. However, these contracts, while serving as transaction protocols, can be compromised by developers with malicious intent, leading to substantial financial losses.
@@
To address this gap in detecting intent in smart contracts, we propose SmartIntentNN, an automated deep learning-based tool designed for smart contract intent detection. It integrates a Universal Sentence Encoder [2] to generate contextual embeddings [3] of smart contracts, a K-means clustering model [4] to identify and highlight intent-related features, and a bidirectional LSTM (long short-term memory) [5, 6] multi-label classification network to predict intents in smart contracts. Evaluations on a dataset of over 10,000 smart contracts show that SmartIntentNN surpasses all baselines, achieving an F1-score of up to 0.8633.
@@
  • We present the first work on smart contract intent detection, utilizing deep learning models.
  • We have compiled an extensible dataset of over 40,000 smart contracts, labeled with 10 categories of intent.
@@ II-B Input Extraction
Smart contract source code on BSC can be published either as single-file contracts with merged imports or as multiple-file contracts. We consolidate multiple files into a single one.
We remove pragma (Solidity compiler version), import statements, and comments as they do not affect intent expression. For multi-file contracts, import statements become redundant after merging.
@@ III-B Intent Highlight
To identify functions that are significantly distant in spatial distribution from these 19 frequently occurring functions, we initially set the number of clusters $k$ to 19 and then conducted a maximum of 80 iterations of K-means clustering training. To compare document similarities, we compute the cosine distance between their embedding vectors[7][8]. Formula 1 defines the cosine similarity between two functions (A and B), derived from the cosine of $\bm{f^{A}}$ and $\bm{f^{B}}$. We then transform the cosine similarity into cosine distance as defined by Formula 2.
@@
In Formula 3, the feature $\bm{f_{i}}$ in matrix $\bm{X}$ is scaled by the predicted within-cluster distance to generate a new matrix $\bm{X^{\prime}}\in\mathbb{R}^{n\times m}$, where $i\in\{1,2,\dots,n\}$ and $\bm{c_{j}}$ represents the cluster centroid, $j\in\{1,2,\dots,16\}$. Here, $\lambda=0.21$ is the threshold; beyond it, $\bm{f_{i}}$ is scaled by a factor of $\mu=16$, referred to as $\mathrm{H_{16}}$ in Section V. This process amplifies rare function code, highlighting their significant intent contribution.
@@ III-C Multi-label Classification
In this section, we utilize a Deep Neural Network (DNN) model for multi-label binary classification. This model comprises three layers: an input layer, a bidirectional LSTM (BiLSTM) layer, and a multi-label classification output layer. The matrix $\bm{X^{\prime}}$ is fed into the model, which is trained by minimizing 10 combined binary cross-entropy losses corresponding to the 10 intent labels described in Section II.A.
@@ IV Application
We developed SmartIntentNN using Tensorflow.js[10], creating a web-based tool accessible through any browser. Specifically, SmartIntentNN offers two primary functionalities: intent highlight and intent detection.
@@ IV-A Intent Highlight
In Fig. 2, several functions are highlighted with a red background, such as setBotBlacklist and setAutoRebase, which indeed exhibit suspicious intent. These functions may correspond to the intent categories of blacklist and rebase described in Section II.A. Non-highlighted functions mainly include interfaces or libraries, such as those in IPancakeSwapFactory.
@@ V Evaluation
This research is pioneering in the field of intent detection in smart contracts and, therefore, has no prior studies for direct comparison. Consequently, we conducted a self-comparison against several established baselines, including models such as LSTM, BiLSTM, and CNN [12]. Furthermore, we benchmarked our model against popular generative large language models for a more comprehensive evaluation.
@@ VI Conclusion
In this research, we introduce SmartIntentNN, a novel automated tool based on deep learning models, designed to detect developers’ intent in smart contracts. SmartIntentNN incorporates a Universal Sentence Encoder, an intent highlight model grounded in K-means, and a DNN integrated with a BiLSTM layer. Trained on $10,000$ and evaluated on $10,000$ distinct smart contracts, SmartIntentNN achieves an F1-score of $0.8633$.

### v4

@@ Abstract
Smart contracts on the blockchain offer decentralized financial services but often lack robust security measures, leading to significant economic losses. While substantial research has focused on identifying vulnerabilities in smart contracts, a notable gap remains in evaluating the malicious intent behind their development. To address this, we introduce SmartIntentNN (Smart Contract Intent Neural Network), a deep learning-based tool designed to automate the detection of developers’ intent in smart contracts. Our approach integrates a Universal Sentence Encoder for contextual representation of smart contract code, employs a K-means clustering algorithm to highlight intent-related code features, and utilizes a bidirectional LSTM-based multi-label classification network to predict ten distinct categories of unsafe intent. Evaluations on 10,000 real-world smart contracts demonstrate that SmartIntentNN surpasses all baselines, achieving an F1-score of 0.8633.
@@ I Introduction
A smart contract is a type of computer program and transaction protocol, engineered to execute, control, or document legally binding events and actions automatically according to the stipulations of a contract or agreement [1]. Users generally interact with smart contracts by initiating transactions to invoke various functions. From a programming standpoint, current research on smart contract security predominantly focuses on identifying vulnerabilities and bugs. However, these contracts, while serving as transaction protocols, can be compromised by developers with malicious intent, leading to substantial financial losses.
@@
To address the gap in detecting intent in smart contracts, we propose SmartIntentNN, an automated deep learning-based tool designed for smart contract intent detection. It integrates a Universal Sentence Encoder (USE) [2] to generate contextual embeddings [3] of smart contracts, a K-means clustering model [4] to identify and highlight intent-related features, and a bidirectional long short-term memory (BiLSTM) [5, 6] multi-label classification network to predict intents in smart contracts. Trained on 10,000 smart contracts and evaluated on another 10,000 distinct contracts, this tool surpasses all baselines, achieving an F1-score of 0.8633.
@@
  • We present the first work on detecting smart contract development intent using deep learning techniques.
  • We compile an extensive dataset of over 40,000 smart contracts, labeled with ten categories of intent.
@@ II-B Input Extraction
Smart contract source code on BSC can be published either as single-file contracts with merged imports or as multiple-file contracts. We consolidate multiple files into a single document.
We remove pragma (Solidity compiler version specification), import statements, and comments as they do not affect intent expression. For multi-file contracts, import statements become redundant after merging.
@@ III-B Intent Highlight
To identify functions that are significantly distant in spatial distribution from these 19 frequently occurring functions, we initially set the number of clusters $k$ to 19 and then conducted a maximum of 80 iterations of K-means clustering training. To compare document similarities, we compute the cosine distance between their embedding vectors [7, 8]. Formula 1 defines the cosine similarity between two functions (A and B), derived from the cosine of $\bm{f^{A}}$ and $\bm{f^{B}}$. We then transform the cosine similarity into cosine distance as defined by Formula 2.
@@
In Formula 3, the feature $\bm{f_{i}}$ in matrix $\bm{X}$ is scaled by the predicted within-cluster distance to generate a new matrix $\bm{X^{\prime}}\in\mathbb{R}^{n\times m}$, where $i\in\{1,2,\dots,n\}$ and $\bm{c_{j}}$ represents the cluster centroid, $j\in\{1,2,\dots,16\}$. Here, $\lambda=0.21$ is the threshold; beyond it, $\bm{f_{i}}$ is scaled by a factor of $\mu=16$, referred to as $\mathrm{H_{16}}$ in Section V. This process amplifies rare functions, highlighting their significant intent contribution.
@@ III-C Multi-label Classification
In this section, we utilize a DNN model for multi-label binary classification. This model comprises three layers: an input layer, a BiLSTM layer, and a multi-label classification output layer. The matrix $\bm{X^{\prime}}$ is fed into the model, which is trained by minimizing 10 combined binary cross-entropy losses corresponding to the 10 intent labels described in Section II.A.
@@ IV Application
We developed SmartIntentNN using Tensorflow.js [10], creating a web-based tool accessible through any browser. Specifically, SmartIntentNN offers two primary modules: intent highlight and intent detection.
@@ IV-A Intent Highlight
In Fig. 2, several functions are highlighted with a red background, such as setBotBlacklist and setAutoRebase, which indeed exhibit suspicious intent. These functions may correspond to the intent categories of blacklist and rebase described in Section II.A. Non-highlighted functions mainly include interfaces or libraries, such as those in IPancakeSwapPair.
@@ V Evaluation
This research is pioneering in the field of intent detection in smart contracts and, therefore, has no prior studies for direct comparison. Consequently, we conducted a self-comparison against several established baselines, including models such as LSTM, BiLSTM, and CNN [12]. Furthermore, we benchmarked our model against popular generative large language models (LLMs) for a more comprehensive evaluation.
@@ VI Conclusion
In this research, we introduced SmartIntentNN, a novel automated tool utilizing deep learning models to detect developers’ intent in smart contracts. SmartIntentNN incorporates a pre-trained USE model, an intent highlight model based on K-means clustering, and a DNN integrated with a BiLSTM layer. Trained on $10,000$ and evaluated on another $10,000$ distinct smart contracts, SmartIntentNN achieved an F1-score of $0.8633$. The model surpasses all baselines, including the latest generative LLMs.